OpenStudy (anonymous):

An aquarium has a volume of 5797 in3. The aquarium has a height of 17 inches and a length of 2 less inches than 3 times the width. Find the length and width of the aquarium.

OpenStudy (imstuck):

the first thing you need to know to do this is the formula for the volume of a rectangular prism. Yes, you know it?

OpenStudy (imstuck):

Well...I'm ready with the answer and to help you understand how to get to it if you are still out there...

OpenStudy (anonymous):

still stuck :/

OpenStudy (anonymous):

this one is also confusing me? A medical plan requires a patient to pay a $10.00 copay plus 30% of all charges insured for a medical procedure. If the average bill for a patient is between $2500 and $3000, find the range of the original bill.

OpenStudy (imstuck):

The first one is set up like this using the fact that we know the volume and the fact that the length is equal to 3w - 2.

OpenStudy (imstuck):

V = L * W * H 5797= (3W - 2)(W)(17) 5797 = 51W^2 - 34W 51W^2 - 34W - 5797 = 0 Now factor that and solve for W. W = -10.33333, W = 11 We all know that length and width cannot be negative numbers so W = 11, L = 3(11) - 2 = 31 and H = 17 (given). Checking those solutions, we should be able to multiply them all together like the volume formula says and end up 5797. Let's check...

OpenStudy (imstuck):

11 * 17 * 31 = 5797

OpenStudy (anonymous):

so lenth = 33 and width = 11 is the answer?

OpenStudy (anonymous):

length

OpenStudy (imstuck):

The length is actually 31, not 33. But the width is 11. So that answers are L = 31 and W = 11.

OpenStudy (anonymous):

thank you :)

OpenStudy (imstuck):

You're welcome!

OpenStudy (anonymous):

how would you solve this one

OpenStudy (anonymous):

A medical plan requires a patient to pay a $10.00 copay plus 30% of all charges insured for a medical procedure. If the average bill for a patient is between $2500 and $3000, find the range of the original bill.

OpenStudy (anonymous):

all I could figure out was the first one was 8300?

OpenStudy (anonymous):

but idk the 2nd range